# On real forms of Fermat hypersurfaces

### v1

@@ 4. The elliptic curve case
gives an elliptic curve $Y^{2}=X^{3}-432$ with $j$-invariant 00, and this can be identified with $E:=\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$ where $\zeta$ is a primitive $6$-th root of unity, which we fix in this section.
@@
with $i\in\mathbb{Z}$ and $a$ in the fundamental domain (cf. [Sil09]). Note that the real structure $\rho\colon F^{1}_{3}\to F^{1}_{3}$ corresponds to $E\to E$ that maps $z$ to $\bar{z}$ where we identify $z\in\mathbb{C}$ with the image of $z$ by the canonical surjection $\mathbb{C}\to\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$. We also denote it by $\rho\colon E\to E$.
@@
$(\alpha\circ\rho)^{2}(z)$ $=\zeta^{i}(\overline{\zeta^{i}\bar{z}+a})+a$
$=z+(\zeta^{i}\bar{a}+a)$
@@
$=$ $\zeta^{i-2j}\bar{z}+\zeta^{-j}a-\zeta^{j}b+\zeta^{i-j}\bar{b}$
@@
we see that $z\to\zeta^{i}z+a$ is equivalent to $z\to\zeta^{i}z$, so the result follows. ∎
@@
by taking $[1:-1:0]$ as a unit. From this, we immediately get the following result:
@@ 5. Explicit description of real forms of Fermat hypersurfaces
First, we consider the curve case $F^{1}_{d}=(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$.
@@
Assume that $d\geq 3$. Then, the correspondence between cocycles and real forms of $F^{1}_{d}$ is given by the following table.
@@
Next, we consider the real form corresponding to $K^{1,2}$. The invariant ring of the homomorphism $\alpha\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, and $X_{2}$ to $X_{2}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}]$ with $\xi_{2d}^{2}=\xi_{d}$.
In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\mu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
$f(X_{0},X_{1},X_{2})=\mu(f)(\xi_{2d}X_{0},X_{1},X_{2}),$
and $\alpha(f)=f$ if and only if $\mu(f)$ is a real coefficient polynomial, as $\xi_{2d}X_{0}$, $X_{1}$, $X_{2}$ are invariant under $\alpha$.
@@
Finally, we consider the real form corresponding to $K^{1,2}$. The invariant ring of the homomorphism $\beta\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $X_{0}$, $X_{1}$ to $X_{2}$ and $X_{2}$ to $X_{2}$ is $\mathbb{R}[X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})]$.
We can show this by a similar argument as above. In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\nu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
$f(X_{0},X_{1},X_{2})=\nu(f)(X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})),$
and $\beta(f)=f$ if and only if $\nu(f)$ is a real coefficient polynomial.
@@
$(Y_{0}^{d}+(Y_{1}+\sqrt{-1})^{d}+(Y_{1}-\sqrt{-1})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
In the surface case, by using the same argument as above, we get the following result:
@@
Assume $d\geq 3,\neq 4$. Then, the correspondence between cocycles and real forms of $F^{2}_{d}$ is given by the following table.
@@
$K^{0,4}_{2}$ (or $H^{4}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
@@
$K^{0,2}_{2}$ (or $H^{2}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
@@
$K^{0,0}_{2}$ (or $H^{0}_{4}$) $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
In the same way, we can extend it to the general case. In particular, all real forms of $F_{d}^{n}$ are the $n$-dimensional hypersurface of degree $d$.
@@
(2) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $x_{n+1}\neq 0$. Therefore, we can define a continuous map $Z\to S^{n}$ as follows:
$[x_{0}:\cdots:x_{n}:1]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots+x_{n}^{2}}}(x_{0},\ldots,x_{n}).$
@@
(3) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{0},\ldots,x_{i})\neq(0,\ldots,0)$, $(x_{i+1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Thus, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto([x_{0}:\cdots:x_{i}],[x_{i+1}:\cdots:x_{n+1}]).$
@@
(4) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Therefore, we can define a continuous function $Z\to\mathbb{P}^{n}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{1}:\cdots:x_{n+1}].$
@@
(5) First, note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{n},x_{n+1})\neq(0,0)$. Therefore, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{n}:x_{n+1}].$
@@
for $10\leq m\leq\frac{d}{2}-$ 1.
In fact, $[x_{n}:x_{n+1}]\in f(Z)$ if and only if $(x_{n}+\sqrt{-1}x_{n+1})^{d}+(x_{n}-\sqrt{-1}x_{n+1})^{d}\leq 0$, and

### v2

@@ 4. The elliptic curve case
gives an elliptic curve $Y^{2}=X^{3}-432$ with $j$-invariant 0, and this can be identified with $E:=\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$ where $\zeta$ is a primitive $6$-th root of unity, which we fix in this section.
@@
with $i\in\mathbb{Z}$ and $a$ in the fundamental domain (cf. [Sil09]). Note that the real structure $\rho\colon F^{1}_{3}\to F^{1}_{3}$ corresponds to $E\to E$ that maps $z$ to $\overline{z}$ where we identify $z\in\mathbb{C}$ with the image of $z$ by the canonical surjection $\mathbb{C}\to\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$. We also denote it by $\rho\colon E\to E$.
@@
$(\alpha\circ\rho)^{2}(z)$ $=\zeta^{i}(\overline{\zeta^{i}\overline{z}+a})+a$
$=z+(\zeta^{i}\overline{a}+a)$
@@
$=$ $\zeta^{i-2j}\overline{z}+\zeta^{-j}a-\zeta^{j}b+\zeta^{i-j}\overline{b}$
@@
we see that $z\to\zeta^{i}z+a$ is equivalent to $z\to\zeta^{i}z$, so, by setting $a=b=0$ in the above calculation, we get the result. ∎
@@
by taking [1:1:0][1:-1:0] as a unit. From this, we immediately get the following result:
@@
5. Degree $2$ cases
(This section is based on the advice from Professor János Kollár. )
In this section, we consider the number and explicit descriptions of the real forms of $X_{n}:=F^{n}_{2}=\left(X_{0}^{2}+\cdots+X_{n+1}^{2}=0\right)\subset\mathbb{P}^{n+1}$.
Let $Q_{r,s}\subset\mathbb{P}^{r+s-1}$ be a quadratic hypersurface defined as follows:
$Q_{r,s}:=\left(X_{0}^{2}+\cdots+X_{r-1}^{2}-X_{r}^{2}-\cdots X_{r+s-1}^{2}=0\right)\subset\mathbb{P}^{r+s-1}.$
Then, $F^{1}_{2}$ is isomorphic to $\mathbb{P}^{1}$, and has exactly two real forms $Q_{2,0}$ and $Q_{1,1}$ up to isomorphism. Also, it is known that $F^{2}_{2}$ has exactly four real forms $Q_{4,0}$, $Q_{3,1}$, $Q_{2,2}$, and $Q_{3,0}\times Q_{2,1}$ up to isomorphisms (cf. [Rus02]).
Therefore, we consider the case where $n\geq 3$.
Lemma 5.1.
When $n\geq 3$, any real forms of $X_{n}$ can be embedded into real forms of $\mathbb{P}^{n+1}$.
Proof.
Let $H$ be a hyperplane section of $X_{n}$. Then, $|H|$ induces an embedding $\varphi_{|H|}\colon X_{n}\hookrightarrow\mathbb{P}^{n+1}$.
Considering the automorphism group of $X_{n}$, we see that $X_{n}$ is also a smooth hypersurface of degree $2$ in $\mathbb{P}^{n+1}$ in this embedding. Then, every real structure of $X_{n}$ preserves $|H|$ as every automorphism of $X_{n}$ is linear. Therefore, by considering the quotient by a real structure, we see that every real form of $X_{n}$ can be embedded into some real form of $\mathbb{P}^{n+1}$. ∎
Theorem 5.2.
When $n$ is odd, any real forms of $F^{2}_{n}$ are isomorphic to $Q_{r,s}$ for some rrss with $r\geq s,r+s=n+2$. When $n$ is even, any real forms of $F^{2}_{n}$ are isomorphic to $Q_{r,s}$ for some rrss with $r\geq s,r+s=n+2$, or the real form corresponding to the cocycle $L_{n}$ in $\operatorname{\operatorname{Aut}}X_{n}$.
In particular, when $n$ is even, $F^{2}_{n}$ has exactly $\frac{n}{2}+3$ real forms, and when $n$ is odd, $F^{2}_{n}$ has exactly $\frac{n+1}{2}+1$ real forms.
Proof.
By [Ben16], $P^{n+1}$ has exactly two real forms up to isomorphisms when $n$ is even, which correspond to the following two real structures:
$\tilde{\rho}$ $\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1};\ [z_{0}:\cdots:z_{n+1}]\mapsto\left[\overline{z_{0}}:\cdots:\overline{z_{n+1}}\right]$
$\tilde{\rho^{\prime}}$ $\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1};\ [z_{0}:\cdots:z_{n+1}]\mapsto\left[-\overline{z_{1}}:\overline{z_{0}}:\cdots:\overline{z_{n+1}}:-\overline{z_{n}}\right].$
Here, we denote by $\widetilde{\mathbb{P}^{n+1}}$ the real form of $\mathbb{P}^{n+1}$ defined by the real structure $\tilde{\rho^{\prime}}$. Also, when $n$ is odd, $\mathbb{P}^{n+1}$ has exactly one real form up to isomorphisms, which corresponds to above $\tilde{\rho}$.
Therefore, by Lemma 5.1, we enough to consider the case where a real form of $F^{2}_{n}$ is embedded into these two real forms of $\mathbb{P}^{n+1}$.
First, suppose a real form $Y$ of $X_{n}$ is embedded into the real form of $\mathbb{P}^{n+1}$ corresponding to the real strucure $\tilde{\rho}$, i.e., $\mathbb{P}^{n+1}_{\mathbb{R}}$. Then, $Y$ is a hypersurface of degree $2$ of $\mathbb{P}^{n+1}_{\mathbb{R}}$ by the proof of Lemma 5.1. As $F^{n}_{2}$ is defined by a non-degenerated quadratic form, $Y$ is isomorphic to $Q_{r,s}$ for some $r,s$ with $r\geq s$ by considering a diagonalization of symmetric matrix.
Second, suppose a real form $Y$ of $X_{n}$ is embedded into the real form of $\mathbb{P}^{n+1}$ corresponding to the real strucure $\tilde{\rho^{\prime}}$, i.e., $\widetilde{\mathbb{P}^{n+1}}$. Then, $\widetilde{\mathbb{P}^{n+1}}$ contains $\widetilde{\mathbb{P}^{1}}$, and as $F^{n}_{2}$ is defined by a non-degenerated quadratic form, we can take $L\subset\widetilde{\mathbb{P}^{n+1}}$ which is isomorphic to $\widetilde{\mathbb{P}^{1}}$ so that it is not contained in $Y$. We may assume that $L=\left(X_{2}=\cdots=X_{n+1}=0\right)\subset\widetilde{\mathbb{P}^{n+1}}$. Also, we may assume that $Y\cap L=\left(X_{0}^{2}+X_{1}^{2}=0\right)\subset L$ as hypersurfaces in $\widetilde{\mathbb{P}^{1}}$ defined by $\left(X_{0}^{2}+X_{1}^{2}=0\right)$ and $\left(X_{0}^{2}-X_{1}^{2}=0\right)$ is isomorphic. Then, by considering orthogonal complement, we can set the coordinate of $\widetilde{\mathbb{P}^{n+1}}$ so that $Y$ is defined by
$X_{0}^{2}+X_{1}^{2}+q\left(X_{2},\ldots,X_{n+1}\right)=0$
where $q$ is a non-degenerated quadratic form. Therefore, by recurring this procedure, we get the unique real form
$\left(X_{0}^{2}+\cdots X_{n+1}^{2}=0\right)\subset\widetilde{\mathbb{P}^{n+1}}$
of $X_{n}$ embedded in $\widetilde{\mathbb{P}^{n+1}}$, and it is clear that this real form correspond the cocycle $L_{n}$ in $\operatorname{\operatorname{Aut}}X_{n}$.
By considering the real loci of $Q_{r,s}$, we see that $Q_{r,s}$’s with $r\geq s$ are not isomorphic to each other. Also, a real form $Q_{r,s}$ of $X_{n}$ corresponds to a cocycle $K^{s,r}_{2}$, and when $n$ is even, this cannot be equivalent to $L_{n}$. In fact, as matrices,
$\begin{pmatrix}-I_{s}&\\ &I_{r}\end{pmatrix}^{2}=I_{r+s}$
whereas
$\begin{pmatrix}0&1&&&\\ -1&0&&&\\ &&\ddots&&\\ &&&0&1\\ &&&-1&0\end{pmatrix}^{2}=-I_{n+2}.$
Therefore, we see that the number of real forms of $X_{n}$ is $\frac{n}{2}+3$ when $n$ is even, and $\frac{n+1}{2}+1$ when $n$ is odd. ∎
@@ 6. Explicit description of real forms of Fermat hypersurfaces
In this section, we consider the explicit description of real forms corresponding to $H^{r}_{n}$, $K^{s,t}_{n}$, and $L_{n}$. We omit subindexes of $H^{r}_{n}$ or $K^{s,t}_{n}$ if it is clear from the context.
First, we consider the real form corresponding to cocycles $L_{n}$. As we see in Theorem 3.5, Theorem 3.6, and Theorem 4.2, cocycles of this type appear only when $n$ and $d$ are even. Let $n=2k$. This is a real form induced by the real form $\widetilde{\mathbb{P}^{n+1}}$ of $\mathbb{P}^{n+1}$, i.e., corresponding to the real structure $\tilde{\rho^{\prime}}$.
Therefore, the quotient $Z^{2k}_{d}$ of $F_{d,\mathbb{C}}^{2k}\subset\mathbb{P}_{\mathbb{C}}^{2k+1}$ with even $d$ by the restriction of $\rho^{\prime}$ is the real form corresponding to $L_{2k}$. Note that this real form can be embedded into $\widetilde{\mathbb{P}^{n+1}}$.
Next, we consider the real form corresponding to cocycles $K$ or $H$. When considering these real forms, the following lemma is useful:
Lemma 6.1.
Let $\xi_{d}$ be a primitive $d$-th root of unity. Then,
  1. (1)
    The invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{1}\colon\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X$ to $\xi_{d}X$ and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}\left[\xi_{2d}X\right]$, with $\xi_{2d}$ satisfying $\xi_{2d}^{2}=\xi_{d}$.
  2. (2)
    The invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{2}\colon\mathbb{C}[X_{1},X_{2}]\to\mathbb{C}[X_{1},X_{2}]$ that sends $X_{1}$ to $X_{2}$, $X_{2}$ to $X_{1}$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}\left[X_{1}+X_{2},\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)\right]$.
Proof.
(1) For all $f\in\mathbb{C}[X]$, there is a unique polynomial $\mu(f)\in\mathbb{C}[X]$ such that
$f(X)=\mu\left(\xi_{2d}X\right).$
Then, as $\xi_{2d}X$ is invariant under $\alpha_{1}$, we see that $\alpha_{1}\left(f\right)=f$ if and only if $\mu(f)$ is a polynomial of real coefficients. This shows that the invariant ring with $\alpha_{1}$ is $\mathbb{R}\left[\xi_{2d}\right]$.
(2) For all $g\in\mathbb{C}[X_{1},X_{2}]$, there is a unique polynomial $\nu(g)\in\mathbb{C}[X_{1},X_{2}]$ such that
$g(X_{1},X_{2})=\nu(g)\left(X_{1}+X_{2},\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)\right).$
Then, as $X_{1}+X_{2}$, $\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)$ is invariant under $\alpha_{2}$, we see that $\alpha_{2}(g)=g$ if and only if $\nu(g)$ is a polynomial with real coefficients. This shows that the invariant ring of $\alpha_{2}$ is $\mathbb{R}\left[X_{1}+X_{2},\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)\right]$. ∎
Then, we consider the curve case $F^{1}_{d}=(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$.
@@
Assume that $d\geq 3$. Then, the correspondence between cocycles and real forms of $F^{1}_{d}$ is given in the following table.
@@
Next, we consider the real form corresponding to $K^{1,2}$. By Lemma 6.1, the invariant ring of the homomorphism $\alpha\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\overline{c}$, $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, and $X_{2}$ to $X_{2}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}]$ with $\xi_{2d}^{2}=\xi_{d}$.
@@
Finally, we consider the real form corresponding to $K^{1,2}$. By Lemma 6.1, the invariant ring of the homomorphism $\beta\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\overline{c}$, $X_{0}$ to $X_{0}$, $X_{1}$ to $X_{2}$ and $X_{2}$ to $X_{1}$ is $\mathbb{R}[X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})]$.
@@
$(Y_{0}^{d}+(Y_{1}+\sqrt{-1}Y_{2})^{d}+(Y_{1}-\sqrt{-1}Y_{2})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
In the surface case, by using a similar argument as above, we get the following result:
@@
Assume $d\geq 3,\neq 4$. Then, the correspondence between cocycles and real forms of $F^{2}_{d}$ with odd $d$ is given in the following table:
@@
$H^{4}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$H^{2}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$H^{0}_{4}$ $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
Also, the correspondence between cocycles and real forms of $F^{2}_{d}$ with even $d$ is given in the following table:
cocycle real form
$K^{0,4}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
@@
$K^{0,2}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
@@
$K^{0,0}_{2}$ $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$L_{2}$ $Z_{d}^{2}$
Proof.
The real form corresponding to $L_{2}$ is already described above. First, we consider the real form corresponding to $K_{2}^{k,l}$ with $k+l=4$. The real form corresponding to $K_{2}^{0,4}$ is clear.
By Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{1}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{0}$ to $\xi_{d}X_{0}$, $X_{i}$ to $X_{i}$ for $i=1,2,3$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2},X_{3}]$. Therefore, by setting
$Y_{0}:=\xi_{2d}X_{0},\ Y_{i}:=X_{i}\ (i=1,2,3),$
we see that the real form corresponding to $K_{2}^{1,3}$ is
$\left(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
Also, by Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{2}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{i}$ to $\xi_{d}X_{i}$ for $i=0,1$, $X_{j}$ to $X_{j}$ for $j=2,3$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[\xi_{2d}X_{0},\xi_{2d}X_{1},X_{2},X_{3}]$. Therefore, by setting
$Y_{i}:=\xi_{2d}X_{i}\ (i=0,1),\ Y_{j}:=X_{j}\ (j=2,3),$
we see that the real form corresponding to $K_{2}^{2,2}$ is
$\left(-Y_{0}^{d}-Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
Next, we consider the real form corresponding to $K_{2}^{k,l}$ with $k+l=2$.
By Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\beta_{1}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{i}$ to $X_{i}$ for $i=0,1$, $X_{2}$ to $X_{3}$, $X_{3}$ to $X_{2}$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[X_{0},X_{1},X_{2}+X_{3},\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)]$. Therefore, by setting
$Y_{i}:=X_{i}\ (i=0,1),\ {Y^{\prime}}_{2}:=X_{2}+X_{3},{Y^{\prime}}_{3}:=\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)$
and also, by setting
$Y_{2}:=\frac{{Y^{\prime}}_{2}}{2},\ Y_{3}:=\frac{{Y^{\prime}}_{3}-{Y^{\prime}}_{2}}{2},$
we see that the real form corresponding to $K_{2}^{1,3}$ is
$\left(Y_{0}^{d}+Y_{1}^{d}+\left(Y_{2}+\sqrt{-1}Y_{3}\right)^{d}+\left(Y_{2}-\sqrt{-1}Y_{3}\right)^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
Also, by Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\beta_{2}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, $X_{2}$ to $X_{3}$, $X_{3}$ to $X_{2}$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}+X_{3},\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)]$. Therefore, by setting
$Y_{0}:=\xi_{2d}X_{0},\ Y_{1}:=X_{1},\ {Y^{\prime}}_{2}:=X_{2}+X_{3},{Y^{\prime}}_{3}:=\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)$
and also, by setting
$Y_{2}:=\frac{{Y^{\prime}}_{2}}{2},\ Y_{3}:=\frac{{Y^{\prime}}_{3}-{Y^{\prime}}_{2}}{2},$
we see that the real form corresponding to $K_{2}^{1,3}$ is
$\left(-Y_{0}^{d}+Y_{1}^{d}+\left(Y_{2}+\sqrt{-1}Y_{3}\right)^{d}+\left(Y_{2}-\sqrt{-1}Y_{3}\right)^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
Finally, we consider the real form corresponding to $K_{2}^{0,0}$. By Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\gamma\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{i}$ to $X_{i+1}$, $X_{i+1}$ to $X_{i}$ for $i=0,2$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[X_{0}+X_{1},\left(1-\sqrt{-1}\right)\left(X_{0}-\sqrt{-1}X_{1}\right),X_{2}+X_{3},\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)]$. Therefore, by setting
${Y^{\prime}}_{i}:=X_{i}+X_{i+1},{Y^{\prime}}_{i+1}:=\left(1-\sqrt{-1}\right)\left(X_{i}-\sqrt{-1}X_{i+1}\right)\ (i=0,2)$
and also, by setting
$Y_{i}:=\frac{{Y^{\prime}}_{i}}{2},\ Y_{i+1}:=\frac{{Y^{\prime}}_{i+1}-{Y^{\prime}}_{i}}{2}\ (i=0,2),$
we see that the corresponding real form is
$\left(\left(Y_{0}+\sqrt{-1}Y_{1}\right)^{d}+\left(Y_{0}-\sqrt{-1}Y_{1}\right)^{d}+\left(Y_{2}+\sqrt{-1}Y_{3}\right)^{d}+\left(Y_{2}-\sqrt{-1}Y_{3}\right)^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
In the same way, we can extend it to the general case. In particular, all real forms of $F_{d}^{n}$ can be described as the $n$-dimensional hypersurface of degree $d$ except for the real form corresponding to $L_{n}$.
@@
(1) As the restriction of $\rho^{\prime}$ to $F^{2k}_{d,\mathbb{C}}$ has no fixed real points, the real locus of $Z^{2k}_{d}$ is $\emptyset$.
@@
(3) Note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $x_{n+1}\neq 0$. Therefore, we can define a continuous map $Z\to S^{n}$ as follows:
$x_{0}:\cdots:x_{n}:1]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots+x_{n}^{2}}}(x_{0},\ldots,x_{n}).$
@@
(4) Note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{0},\ldots,x_{i})\neq(0,\ldots,0)$, $(x_{i+1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Thus, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$ as follows:
$x_{0}:\cdots:x_{n+1}]\mapsto([x_{0}:\cdots:x_{i}],[x_{i+1}:\cdots:x_{n+1}]).$
@@
(5) Note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Therefore, we can define a continuous function $Z\to\mathbb{P}^{n}(\mathbb{R})$ as follows:
$x_{0}:\cdots:x_{n+1}]\mapsto[x_{1}:\cdots:x_{n+1}].$
@@
(6) First, note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{n},x_{n+1})\neq(0,0)$. Therefore, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})$ as follows:
$x_{0}:\cdots:x_{n+1}]\mapsto[x_{n}:x_{n+1}].$
@@
for $0\leq m\leq\frac{d}{2}-1$.
In fact, $x_{n}:x_{n+1}]\in f(Z)$ if and only if $(x_{n}+\sqrt{-1}x_{n+1})^{d}+(x_{n}-\sqrt{-1}x_{n+1})^{d}\leq 0$, and